OpenStudy (safyousuff97):

Select the equation of the line that passes through the point (5, 7) and is perpendicular to the line x = 4. y = 4 x = 5 y = 7 x = 7

OpenStudy (safyousuff97):

those r the answer choices ^^

OpenStudy (anonymous):

y=7- is horizontal line, therefore parallel to x=4, and passes through the point.

OpenStudy (mathmale):

First, please find the slope of a line that is perpendicular to the line x = 4. Next, use the point-slope formula y2 - y1 = m(x2-x1) to find the equation of the new line.